The ages of all kindergarten children are normally distributed with
mean 4 years and standard deviation 6 months.
What is the probability that a randomly chosen kindergarten child
will be older than 5 years, 4 months old?
Use Table A-2 and write your answer as a decimal with 4 decimal places.

Answer by stanbon(75887)

The ages of all kindergarten children are normally distributed with
mean 4 years and standard deviation 6 months.
----------------------------------------
What is the probability that a randomly chosen kindergarten child
will be older than 5 years, 4 months old?
-----
Note: 5 yrs 4 months = 5 1/3 = 16/3 yrs
-------
z(16/3) = [(16/3)-(12/3))/(18/3) = (4/3)/(18/3) = 2/9
-----------
P(x > 5yrs 4 mths) = P(z > 2/9) = normalcdf[(2/9),100] = 0.4121
=====================
Cheers,
Stan H.
=====================